OpenStudy (anonymous):

Which is the equation of a line that passes through the point (3, -4) and has a slope of 2. Select one: a. y = 2x - 10 b. y = 2x - 2 c. y = 2x - 4 d. y = 2x - 11

OpenStudy (campbell_st):

well the slope intercept form of the line is y = mx + b m = slope and b = y-intercept you know m = 2 so the equation is y = 2x + b to find b, substitute the point x = 3 and y = -4 hope it helps

OpenStudy (haseeb96):

first find b from the above explanation of @campbell_st

OpenStudy (anonymous):

@campbell_st what do you mean plug it in for x? like y=2(3)+b is y=6+b or y=8+b what do i do after

OpenStudy (campbell_st):

almost y = -4 and x = 3 so its \[-4 = 2 \times 3 + b\] now solve for b

OpenStudy (haseeb96):

first find b like this y=2x+b..........eq#1 -4 = 2(3) + b b=-10 now put b=-10 into eq#1 and you will get y= 2x-10

OpenStudy (campbell_st):

there you go... someone always gives the answer

OpenStudy (haseeb96):

@campbell_st sorry if you dont like it but medal is your

OpenStudy (haseeb96):

@AWillie98 did you get it how we solve it ?

OpenStudy (anonymous):

i didn't want @Haseeb96 to give me the answer @campbell_st ill never learn it that way